Using divisibility tests, determine which of the following numbers are divisible by 11: 
 7138965 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of digits at odd places = 5 + 9 + 3 + 7 = 24
Sum of digits at even places = 6 + 8 + 1 = 15 Difference = 24–15 = 9
Since, the difference between sum of digits at odd places and sum of digits at even places is 9 which is not divisible by 11.
Hence, 7138965 is not divisible by 11

​Thanks